### v1

# Efficient absorbing boundary conditions for conservation laws: Application to the piston problem of gas dynamics

Abstract
In this work, we address the imposition of outflow boundary conditions for one-dimensional conservation laws. While a highly accurate numerical solution can be obtained in the interior of the domain, boundary discretization can lead to unphysical reflections. We investigate and implement some classes of relaxation methods and far-field operators to deal with this problem without significantly increasing the size of the computational domain. Relations are established between these techniques, and extensions of them are explored. In particular, we introduce a simple and robust relaxation method with a matrix-valued weight function that selectively absorbs outgoing waves. As a challenging model problem, we consider the Lagrangian formulation of the Euler equations for an isotropic gas with inflow boundary conditions determined by an oscillating piston.
1 Introduction
Hyperbolic systems of partial differential equations (PDEs) are characterized by finite propagation speed of information in their solutions. At a fixed time, and with given bounded and compactly supported initial data, we expect the relevant information of a phenomenon described by a system of conservation laws to be concentrated inside a bounded subset of the domain. Nevertheless, when the physical domain is infinite, increasingly extensive computational domains might be necessary, which can become prohibitively expensive. Even though this has been partially alleviated by the fast development of computer hardware, to accurately capture the long-time behavior of solutions still poses a significant challenge. As a consequence, a variety of absorbing boundary conditions (ABCs) have been proposed to tackle this problem (see, e.g., [8, 13] and references therein).
In this manuscript, we focus on the so-called relaxation and far-field operators methods, briefly outlined below. These are families of ABCs that could, in principle, be applied to any system of hyperbolic PDEs. Both of them rely on the use of a region of the domain where the numerical solution is damped or forced towards a given function. To avoid ambiguity, we will refer to this region as the sponge layer throughout this work (see Figure 1).
Larsen and Dancy [25] introduced the relaxation method to apply ABCs to their discretization of Peregrine’s Boussinesq type equations. The method was first presented as a postprocessing technique where the numerical solution is divided by an increasing set of positive numbers inside the sponge layer (often referred to as relaxation zone in the literature) at each step of the time integration procedure. A Fourier mode-based reflection analysis of a linearization of Peregrine’s equations suggested that the absorbing properties of this technique improve drastically as the ratio between the the length of the sponge layer and the wavelength increases. Due to the versatility and efficiency of this method, it has been widely used for both absorption and generation of waves in water wave modeling, where validation tests often consist of solitary waves with a known characteristic length that can be contained inside the sponge layer [28, 6, 17]. Furthermore, several works that formalize this technique and analyze its non-reflecting properties have been proposed, e.g. [30, 4, 5, 14].
Another popular approach for the development of ABCs is the modification of the system of governing equations, such that damping of outgoing waves is a characteristic behavior of the analytic solution. In their seminal work [16], Israeli and Orszag discuss the addition of reaction and diffusion terms with variable strength inside the sponge layer. They further explore the idea of absorbing waves in just one direction for the wave equation by modifying its second-order scalar and first-order system forms. However, it was later shown by Karni [19] that such modifications unavoidably generate reflections that propagate into the computational domain. Instead, Karni proposes the use of far-field slowing-down and damping operators that, by preserving the eigenstructure of the flux Jacobian, absorb waves in a prescribed direction for systems of hyperbolic PDEs. A generalization of Karni’s slowing-down and damping operators for linear systems was introduced by Appelö and Colonius [1], where high-order viscosity terms are added in the sponge layer.
As a benchmarking application for the ABCs discussed in this work, we consider the one-dimensional Euler equations in their Lagrangian formulation. These equations are used to model the piston problem of gas dynamics, where the solution is known to consist of quasi-periodic nonlinear plane waves with strength proportional to the amplitude of the source [40]. In particular, we study the strongly nonlinear case, where shock formation occurs in the near-field and recurrent interaction between the first and second shocks of the wave train takes place in the far-field, as demonstrated by Inoue and Yano [15]. A central scheme discretization for this problem and its multi-fluid counterpart in Lagrangian coordinates was put forward by Fazio and Russo [7], where the prescription of second-order accurate inflow boundary conditions at the piston, used herein, is described.
The specific contributions of the present work are:
  • a simple relation between the relaxation method and far-field damping operators ABCs for general hyperbolic systems of conservation laws (see Sections 3.1 and 3.3);
  • A generalization of the relaxation method to improve its absorbing properties without significantly increasing its computational cost (see Section 3.2);
  • a new damping operator based on nonlinear characteristic variables (see Section 3.4);
  • a detailed performance comparison of different state-of-the-art ABCs for the Lagrangian formulation of the Euler equations with emphasis on the piston problem of gas dynamics (see Section 5).
This manuscript is organized as follows. In Section 2 some existing sponge layer-based ABCs are reviewed. In Section 3, we introduce extensions of the techniques discussed in Section 2. In Section 4 we recall the compressible Euler equations for an isotropic gas, their Lagrangian formulation, and the piston problem of gas dynamics. In Section 5 we assess the performance of the different ABCs in the context of the piston problem.
The code used for the development of this work is available in the reproducibility repository [31].
Figure 1: Outgoing waves are absorbed in the sponge layer through the application of artificial boundary conditions.
2 Sponge layer-based ABCs
Unless stated otherwise, hereinafter we consider the governing equations to be a general hyperbolic system of PDEs of the form
$\partial_{t}\mathbf{q}+\partial_{x}\mathbf{f}(\mathbf{q})={\bf 0}.$ (1)
Here $\mathbf{q}\in L^{1}([0,T]\times[0,+\infty[,^{N})$, and the Jacobian matrix of $\mathbf{f}$ can be diagonalized as $\mathbf{f}^{\prime}(\mathbf{q})=\mathbf{R}\mathbf{\Lambda}{\mathbf{R}}^{-1}$, where the columns of $\mathbf{R}$ are the right eigenvectors $\{\mathbf{r}_{i}(\mathbf{q})\}_{i=1}^{N}$ of $\mathbf{f}^{\prime}(\mathbf{q})$, and $\boldsymbol{\Lambda}$ is a diagonal matrix containing its real eigenvalues $\{\lambda_{i}(\mathbf{q})\}_{i=1}^{N}$. The linear case, where $\mathbf{f}(\mathbf{q})=\mathbf{A}\mathbf{q}$ for some constant matrix $\mathbf{A}\in^{N\times N}$,
$\partial_{t}\mathbf{q}+\mathbf{A}\partial_{x}\mathbf{q}={\bf 0},$ (2)
will be of relevance for some of the techniques discussed below. Throughout this work, it is assumed that an appropriate discretization technique is available for these systems, and a particular approach is briefly discussed in Section 5.1.
The computational domain and sponge layer will be the intervals $[0,x_{s}[$ and $[x_{s},x_{\infty}]$ respectively, and we suppose there is a far-field steady-state $\overline{\mathbf{q}}\in^{N}$ such that $\mathbf{q}(0,x)=\overline{\mathbf{q}}$ for $x>x_{s}$. Our goal is to impose ABCs in the sponge layer so that perturbations of $\overline{\mathbf{q}}$ in the solution are absorbed before they reach the right boundary $x_{\infty}$, as depicted in Figure 1. Regarding the left boundary, we assume the existence of a function $\mathbf{q}_{BC}(\cdot)$ such that well-posedness is guaranteed with the inflow boundary condition $\mathbf{q}(t,0)=\mathbf{q}_{BC}(t)$.
2.1 Slowing-down and damping far-field operators (SDO)
In the work [19], Karni proposes to impose ABCs for the linear system (2) by solving the modified system
$\partial_{t}\mathbf{q}+\mathbf{S}(x)\partial_{x}\mathbf{q}=-\mathbf{D}(x)\left(\mathbf{q}-\overline{\mathbf{q}}\right),$ (3)
instead. $\mathbf{S}$ and $\mathbf{D}$ are referred to as slowing-down and damping far-field operators respectively, and are given by
$\mathbf{S}(x)=\mathbf{R}{\bf\Lambda_{S}}(x)\mathbf{R}^{-1},\quad\mathbf{D}(x)=\mathbf{R}{\bf\Lambda_{D}}(x)\mathbf{R}^{-1},$ (4)
where ${\bf\Lambda_{S}}(x)$ and ${\bf\Lambda_{D}}(x)$ are diagonal matrices with entries $\{s_{i}(x)\lambda_{i}\}_{i=1}^{N}$ and $\{d_{i}(x)s_{i}(x)\lambda_{i}\}_{i=1}^{N}$.
After writing (3) in terms of the characteristic variables $\mathbf{w}=\mathbf{R}^{-1}\mathbf{q}$, and if $s_{i}$ and $d_{i}$ are constant, we have $N$ decoupled advection-reaction equations
$\partial_{t}w_{i}+s_{i}\lambda_{i}\partial_{x}w_{i}=-d_{i}s_{i}\lambda_{i}(w_{i}-\overline{w}_{i}),$ (5)
with explicit solution
$w_{i}(x,t)=\exp(-d_{i}x)\left[\overset{\circ}{w}_{i}(x-s_{i}\lambda_{i}t)-\overline{w}_{i}\right],$ (6)
where $\overset{\circ}{\mathbf{q}}(x)$ is the initial condition for the homogeneous problem. Thus, the speed of propagation and strength of waves corresponding to the $i$-th characteristic field will be reduced at the rates $s_{i}$ and $d_{i}$ respectively. The rationale behind these modifications is that outgoing waves (corresponding to $\lambda_{j}>0$ for an outlet at the right of the domain) can be both slowed down and damped by choosing $s_{j}<1$ and $d_{j}>0$ inside the sponge layer. This way, outgoing waves neither reach the boundary nor reflect. Furthermore, physical waves that are expected to propagate back into the computational domain remain unperturbed.
Karni shows that by choosing the operators as in Equation (4) there are no reflections at an interface point $x$ such that $s_{j}({x}^{-})\neq s_{j}({x}^{+})$ or $d_{j}({x}^{-})\neq d_{j}({x}^{+})$ at the analytical level. However, at the numerical level, reflections proportional to the jump in the coefficients $s_{j}$ and $d_{j}$ are expected [19]. Therefore, to absorb the $j$-th wave, $s_{j}(x):=\Gamma(x)$ and $d_{j}(x):=d(x)$ should be smooth functions as depicted in Figure 3. Moreover, coarser grids and narrower sponge layers will lead to larger unwanted reflections. Since we are interested in absorbing right-going waves, we can set $s_{j}\equiv 1$, $d_{j}\equiv 0$ for $\lambda_{j}\leq 0$, and $0\leq s_{j}(x)\leq 1$, $d_{j}(x)\geq 0$ for $\lambda_{j}>0$. In the following, we will use the abbreviation SDO to refer to this particular choice.
If all the waves are slowed down and damped at the same rate instead, System (3) becomes
$\partial_{t}\mathbf{q}+s(x)\mathbf{A}\partial_{x}\mathbf{q}=-d(x)s(x)\mathbf{A}\left(\mathbf{q}-\overline{\mathbf{q}}\right).$ (7)
Damping source terms with a similar form have been applied previously to, e.g., the wave equation [16], the linearized Euler equations [9], and the Schrödinger equation [23]. For the sake of comparison, we will also consider (7) and refer to it as scalar slowing-down and damping operators (S-SDO).
Even though the SDO technique was introduced for a linear system, a natural way to use it for nonlinear systems is to compute the far-field terms (4) (now dependent on $\mathbf{q}$) locally [18]. This will be the approach considered in Section 5 for the nonlinear equations (39). Additionally, an alternative extension to the nonlinear setting is proposed in Section 3.4.
2.2 The relaxation method (RM)
The relaxation method is a postprocessing technique where the numerical solution is forced towards a prescribed function after each time step. The process is described in Algorithm 1, where the so-called weight function $\Gamma(x)$ decreases from $1$ to 00 in the sponge layer, as depicted in Figure 3. In the relaxation step, $\mathbf{q}$ is replaced by a convex combination of the numerical solution $\mathbf{q}^{*}$ and some target function, which in this case is the constant far-field state $\overline{\mathbf{q}}$.
$\mathbf{q}^{[n]}$
$\mathbf{q}^{[n+1]}$
$\mathbf{q}^{*}\leftarrow$ Evolve a discretization of (1) over a time step $\Delta t$ with initial data $\mathbf{q}^{[n]}$triangleright Hyperbolic step
$\mathbf{q}^{[n+1]}\leftarrow\Gamma(x)\mathbf{q}^{*}+(1-\Gamma(x))\overline{\mathbf{q}}$triangleright Relaxation step
Algorithm 1 Relaxation method with scalar weight function
It is well known that the performance of this ABC depends on the choice of the weight function [3]. By considering the method as a nodal filtering procedure, Engsig-Karup [5] shows that a necessary condition to preserve the continuity of the modal coefficients through elements is that
$\Gamma(x_{s})=1,\quad\Gamma^{\prime}(x_{s})=\Gamma^{\prime\prime}(x_{s})=\Gamma^{\prime\prime\prime}(x_{s})=...=0,$ (8)
which is similar to the smooth transition requirement on the slowing-down and damping functions proposed by Karni [19]. Based on this observation, in [4] the weight function
$\Gamma_{A}(x)=-2\left(1-\phi(x)\right)^{3}+3\left(1-\phi(x)\right)^{2},$ (9)
is proposed, where
$\phi(x)=\left(\frac{x-x_{s}}{x_{\infty}-x_{s}}\right)\chi_{]x_{s},x_{\infty}]}(x)+\chi_{]x_{\infty},+\infty[}(x).$ (10)
In [30], Mayer et al. propose the weight function
$\Gamma_{B}(x)=1-\left[b\phi(x)^{3}+(1-b)\phi(x)^{6}\right],$ (11)
where $b\in[0,1]$ is a parameter that determines its rate of decrease.
The accuracy of this method also depends on the sponge layer length. A customary rule of thumb to design sponge layers for the relaxation method is the requirement $|x_{\infty}-x_{s}|>2L$, where $L$ is the largest wavelength present in the solution [30, 4, 3]. While this restriction might be inconvenient for problems involving long waves, it is not unique to the relaxation method. For instance, Perić and Abdel-Maksoud [34] propose the same requirement on the sponge layer for their linear and quadratic damping far-field operators.
In order to improve the properties of this method, a popular approach is to define a specialized weight function. Hu et al. [14] choose the weight function at each time step such that the mass absorbed artificially is equal to the mass entering the sponge layer. They report that, with this approach, acceptable results are attained with the looser restriction $|x_{\infty}-x_{s}|>0.4L$ for the 2D Navier-Stokes equations with free surface. In [36], Seng notes that the damping effect of the relaxation method increases dramatically as the time step size is refined. To mitigate this phenomenon, the author proposes a weight function that is updated dynamically depending on the distance between the numerical and target solutions.
A variation of the relaxation method was proposed by Wasistho et al. [39] for the 2D Navier-Stokes equations, where the authors proposed the application of relaxation after each stage of the time integration with a Runge-Kutta scheme. Some numerical results concerning this approach are presented in Section 5.3.
3 Extensions of some sponge layer-based ABCs
3.1 A relation between the RM and the far-field damping operator
By performing some manipulations to the relaxation step in Algorithm 1, we obtain
$\mathbf{q}^{[n+1]}$ $=\Gamma(x)\mathbf{q}^{*}+(1-\Gamma(x))\overline{\mathbf{q}}$ (12a)
$=\mathbf{q}^{*}+\Delta t\frac{1-\Gamma(x)}{\Delta t}(\overline{\mathbf{q}}-{\mathbf{q}})$ (12b)
$=:\mathbf{q}^{*}+\Delta t\,\tilde{\Gamma}(x;\Delta t)(\overline{\mathbf{q}}-\mathbf{q}^{*}),$ (12c)
which is a forward Euler step to approximate the solution $\mathbf{q}(\Delta t)$ to the system
$\partial_{t}\mathbf{q}=\tilde{\Gamma}(x;\Delta t)(\overline{\mathbf{q}}-\mathbf{q}),\quad\mathbf{q}(0)=\mathbf{q}^{*}.$ (13)
Hence, Algorithm 1 can be seen as a discretization of
$\partial_{t}\mathbf{q}+\partial_{x}\mathbf{f}(\mathbf{q})=\tilde{\Gamma}(x;\Delta t)(\overline{\mathbf{q}}-\mathbf{q}),$ (14)
using Lie-Trotter operator splitting.
The right hand side of Equation (14) can be written as
$\tilde{\Gamma}(x;\Delta t)(\overline{\mathbf{q}}-\mathbf{q})$ $=\tilde{\Gamma}(x;\Delta t)\mathbf{R}\mathbf{R}^{-1}(\overline{\mathbf{q}}-\mathbf{q})$ (15a)
$=\mathbf{R}\begin{pmatrix}\frac{\tilde{\Gamma}(x;\Delta t)\lambda_{1}}{\lambda_{1}}&0&\cdots&0\\ 0&\frac{\tilde{\Gamma}(x;\Delta t)\lambda_{2}}{\lambda_{2}}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&\frac{\tilde{\Gamma}(x;\Delta t)\lambda_{N}}{\lambda_{N}}\end{pmatrix}\mathbf{R}^{-1}(\overline{\mathbf{q}}-\mathbf{q})$ (15b)
$=\mathbf{R}{\bf\Lambda_{D}}(x;\Delta t)\mathbf{R}^{-1}(\overline{\mathbf{q}}-\mathbf{q})=\mathbf{D}(x;\Delta t)(\overline{\mathbf{q}}-\mathbf{q}),$ (15c)
using the notation from Section 2.1. It follows that, with the RM, the $i$-th characteristic field is absorbed inside the sponge layer at the damping rate
$d_{i}^{\text{RM}}(x;\Delta t)=\frac{\tilde{\Gamma}(x;\Delta t)}{\lambda_{i}}=\frac{1-\Gamma(x)}{\Delta t\lambda_{i}}.$ (16)
Moreover, the strength of left-going waves ($\lambda_{j}<0$) will be amplified according to Equation (6), since $d_{j}^{RM}\leq 0$. The inversely proportional relation between $d_{i}$ and $\Delta t$ is consistent with the increase in damping severity as the time step size decreases using the RM observed in the literature (see, e.g., [36]).
3.2 The RM with matrix-valued weight function (RM-M)
We propose a simple extension of the RM such that it resembles the directional behavior of Karni’s damping far-field operator. We achieve this by allowing the weight function ${\mathbf{\Gamma}}$ to be matrix-valued and defined in terms of $\mathbf{R}$ (and therefore $\mathbf{q}$ in the nonlinear case). The process is described in Algorithm 2 and justified as follows. Proceeding analogously to the previous section, the relaxation step in Algorithm 2
$\mathbf{q}^{[n+1]}={\bf\Gamma}(x)\mathbf{q}^{*}+(\mathbf{I}-{\bf\Gamma}(x))\overline{\mathbf{q}},$ (17)
can be seen as a forward Euler discretization of
$\partial_{t}\mathbf{q}=\tilde{\bf{\Gamma}}(x,\mathbf{q};\Delta t)(\overline{\mathbf{q}}-\mathbf{q}),$ (18)
where
$\tilde{\bf{\Gamma}}(x,\mathbf{q};\Delta t)=\frac{1}{\Delta t}(\mathbf{I}-{\bf\Gamma}(x))=\mathbf{R}\begin{pmatrix}\frac{1-\Gamma_{1}(x)}{\Delta t}&0&\cdots&0\\ 0&\frac{1-\Gamma_{2}(x)}{\Delta t}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&\frac{1-\Gamma_{N}(x)}{\Delta t}\end{pmatrix}\mathbf{R}^{-1}.$ (19)
Thereby, up to first order of accuracy in time, we approximate a system of the form (3), where the desired outgoing wave is damped at the rate
$d_{i}^{\text{RM-M}}(x)=\frac{1-\Gamma_{i}(x)}{\Delta t\lambda_{i}},$ (20)
while the other waves remain unperturbed. The relaxation method of Section 2.2 is a particular case of this formulation, where all the waves are modified at the non-zero rate (20) inside the sponge layer.
If $\mathbf{R}$ is independent of $\mathbf{q}$, the matrix-valued weight function ${\mathbf{\Gamma}}$ does not have to be computed at each time step in Algorithm 2. Additionally, if the far-field state $\overline{\mathbf{q}}$ is constant, we can define the translated function $\mathbf{v}=\mathbf{q}-\overline{\mathbf{q}}$ that satisfies the conservation law $\partial_{t}\mathbf{v}+\partial_{x}\mathbf{f}(\mathbf{v}+\overline{\mathbf{q}})=0$. This system may be solved with the far-field state $\bf 0$ and $\mathbf{q}$ recovered afterward. In this case, we have
$||\mathbf{v}^{[n+1]}||_{2}\leq||{\mathbf{\Gamma}}\mathbf{v}^{*}||_{2}\leq||{\mathbf{\Gamma}}||_{2}||\mathbf{v}^{*}||_{2}\leq||\mathbf{v}^{*}||_{2},$ (21)
thus the stability of the relaxation step is guaranteed.
$\mathbf{q}^{[n]}$
$\mathbf{q}^{[n+1]}$
$\mathbf{q}^{*}\leftarrow$ Evolve a discretization of (1) over a time step $\Delta t$ with initial data $\mathbf{q}^{[n]}$triangleright Hyperbolic step
for each characteristic field $(\lambda_{i},\mathbf{r}_{i})$ do
     if $i$-th wave is being damped then
         $\Gamma_{i}(x)\leftarrow\Gamma(x)$
     else
         $\Gamma_{i}(x)\leftarrow 1$
     end if
end for
${\bf\Lambda}_{\mathbf{\Gamma}}(x)\leftarrow$diagonal matrix-valued function with entries $\Gamma_{i}(x)$
${\bf{\Gamma}}(x)\leftarrow\mathbf{R}{\bf\Lambda}_{\mathbf{\Gamma}}(x)\mathbf{R}^{-1}$ where $\mathbf{R}$ contains the right eigenvectors of $\mathbf{f}^{\prime}(\mathbf{q})$
$\mathbf{q}^{[n+1]}\leftarrow{\bf\Gamma}(x)\mathbf{q}^{*}+(\mathbf{I}-{\bf\Gamma}(x))\overline{\mathbf{q}}$triangleright Relaxation step
Algorithm 2 Directional relaxation method with matrix-valued weight function
3.3 On attaining higher order of accuracy in time with the RM
Let us consider Algorithm 2 for the linear problem (2), where the matrix-valued weight function $\mathbf{\Gamma}$ is independent of $\mathbf{q}$. Then, the relaxation step (17) could be written as a high-order discretization of a system $\partial_{t}\mathbf{q}=\mathcal{R}(g(\mathbf{\Gamma}(x),\Delta t))\,(\overline{\mathbf{q}}-\mathbf{q})$, where $\mathcal{R}$ is the stability function of some Runge-Kutta scheme [10]. Moreover, we can interpret it as the exact solution to the problem
$\partial_{t}\mathbf{q}=\mathbf{W}(x;\Delta t)(\overline{\mathbf{q}}-\mathbf{q}),\quad\mathbf{q}(0)=\mathbf{q}^{*},$ (22)
where $\mathbf{W}$ satisfies
${\bf\Gamma}(x;\Delta t)=\exp({-\Delta t\mathbf{W}(x;\Delta t)}).$ (23)
Since ${\bf\Gamma}=\mathbf{R}{\bf\Lambda}_{\mathbf{\Gamma}}\mathbf{R}^{-1}$, then $\mathbf{W}=\mathbf{R}{\bf\Lambda_{W}}\mathbf{R}^{-1}$, where
$w_{i}(x)=\frac{-\ln\left(\Gamma_{i}(x)\right)}{\Delta t},$ (24)
with $\Gamma_{i}$ and $w_{i}$ denoting the main diagonal entries of the matrices ${\bf\Lambda}_{\mathbf{\Gamma}}$ and ${\bf\Lambda_{W}}$ respectively. Thus, the $i$-th wave is damped at the rate
$d_{i}^{\mathbf{W}}(x)=\frac{w_{i}(x)}{\lambda_{i}}=\frac{-\ln\left(\Gamma_{i}(x)\right)}{\Delta t\lambda_{i}},$ (25)
and will be completely absorbed as $\Gamma_{i}\to 0$, i.e., at the boundary of the sponge layer $x_{\infty}$. This further justifies the approaches proposed by Seng [36] and Wasistho et al. [39] where, through different arguments, $\Gamma$ is substituted by $\Gamma^{C\Delta t}$ to reduce the damping sensitivity to the time step size.
As a consequence of the Lie-Trotter splitting, a first-order error in time can be expected when approximating a solution to the system
$\partial_{t}\mathbf{q}+\partial_{x}\mathbf{f}(\mathbf{q})=\mathbf{W}(x;\Delta t)(\overline{\mathbf{q}}-\mathbf{q}),$ (26)
and a natural approach to get a higher-order accurate in time approximate solution to (26) would be to use the RM in the context of high-order operator splitting. For instance, we could use Strang splitting as described in Algorithm 3. However, in this setting, both splitting techniques are equivalent up to the choice of the scalar weight functions $\Gamma_{i}$. This can be seen as follows: Assuming that the solution is initialized as $\overline{\mathbf{q}}$ inside the sponge layer, the first relaxation step at $t^{[0]}=0$ does not alter the solution. Similarly, the second relaxation step in Algorithm 3 at the final time does not affect the solution in the computational domain, since it is local and no hyperbolic step is taken afterward. Finally, the following relation holds between $\mathbf{q}^{**}$ (computed at time $t^{[n]}$) and $\mathbf{q}^{*}$ (computed at time $t^{[n+1]}$)
$\mathbf{q}^{*}=\mathbf{\Gamma}\mathbf{q}^{[n+1]}+(\mathbf{I}-\mathbf{\Gamma})\overline{\mathbf{q}}=\mathbf{\Gamma}(\mathbf{\Gamma}\mathbf{q}^{**}+(\mathbf{I}-\mathbf{\Gamma})\overline{\mathbf{q}})+(\mathbf{I}-\mathbf{\Gamma})\overline{\mathbf{q}}=\mathbf{\Gamma}^{2}\mathbf{q}^{**}+(\mathbf{I}-\mathbf{\Gamma}^{2})\overline{\mathbf{q}}.$ (27)
The latter amounts to taking a single relaxation step, i.e., using Algorithm 2, with the modified scalar weight functions $\Gamma_{i}^{2}$. Therefore, in the linear case, as long as the scalar weight functions $\Gamma_{i}$ are tuned appropriately for the application at hand, a second-order in-time approximation of a system of the form (26) is obtained with Algorithm 2.
On the other hand, since $\mathbf{\Gamma}$ depends on $\mathbf{q}$ in the nonlinear case, the last observation is no longer valid, and Algorithms 2 and 3 are not equivalent a priori. The effect of applying the RM as in Algorithm 3 in the nonlinear setting is evaluated numerically in Section 5.3.
$\mathbf{q}^{[n]}$
$\mathbf{q}^{[n+1]}$
Compute the weight function ${\bf\Gamma}(x)$ as in Algorithm 2
[…]
$\mathbf{q}^{**}\leftarrow$ Evolve a discretization of (1) over a time step $\Delta t$ with initial data $\mathbf{q}^{*}$triangleright Hyperbolic step
[…]
Algorithm 3 Relaxation method with Strang splitting
3.4 Nonlinear damping far-field operator (NDO)
We propose an alternative extension of the damping operator from Section 2.1 to hyperbolic systems of the form (1) with a general nonlinear flux function. Here, the change of variables $\mathbf{w}=\mathbf{R}^{-1}\mathbf{q}$ no longer allows us to decompose a nonlinear system of the form (1) into a system of scalar evolution equations. Instead, we consider the nonlinear characteristic variables $\boldsymbol{\alpha}=\mathbf{R}^{-1}\partial_{x}\mathbf{q}$ which, after some manipulations to System (1) (see [27, Ch.1]), yield
$\partial_{t}\alpha_{i}+\partial_{x}\left(\lambda_{i}(\mathbf{q})\alpha_{i}\right)=\mathcal{I}(\mathbf{q},\boldsymbol{\alpha}),\quad 1\leq i\leq N,$ (28)
where $\mathcal{I}(\mathbf{q},\boldsymbol{\alpha})$ denotes the nonlinear interactions between the $N$ characteristic fields. Although this is not a decoupled system of equations, we can add a damping term at the right-hand side, resembling the modification done in (5) for the linear equations (2), to obtain
$\partial_{t}\alpha_{i}+\partial_{x}\left(\lambda_{i}(\mathbf{q})\alpha_{i}\right)=\mathcal{I}(\mathbf{q},\boldsymbol{\alpha})-d_{i}(x)\lambda_{i}(\mathbf{q})\alpha_{i},\quad 1\leq i\leq N.$ (29)
After multiplying this system from the left by $\mathbf{R}$ and integrating from $x$ to +$\infty$+ ∞ (inverting the process carried out to obtain (28)), we get
$\partial_{t}\mathbf{q}(t,x)+\partial_{x}\mathbf{f}(\mathbf{q}(t,x))$ $=\int_{x}^{x_{\infty}}\mathbf{R}\boldsymbol{\Lambda}_{\mathbf{D}}(z){\mathbf{R}}^{-1}\partial_{z}\mathbf{q}(t,z)\,dz$ (30a)
$=\int_{x}^{x_{\infty}}\mathbf{D}(z,\mathbf{q})\partial_{z}\mathbf{q}(t,z)\,dz,$ (30b)
where the artificial boundary at the end of the sponge layer is enforced through the assumption $\mathbf{q}(t,x)=\overline{\mathbf{q}}$ for $x\geq x_{\infty}$. If the damping rate functions $d_{i}$ are defined as in Section 2.1, the term (30b) vanishes outside of the sponge layer. Moreover, if the flux is linear and each $d_{i}$ is constant, the damping far-field operator of Section 2.1 is recovered.
Some challenges that arise when dealing with the nonlinear damping term (30b) are its non-local nature and the presence of a nonconservative product. This term is significantly simplified if one can write $\mathbf{D}(z,\mathbf{q})=\partial_{\mathbf{q}}\mathbf{h}(z,\mathbf{q})$ for some function $\mathbf{h}\in C^{1}(\times^{N},^{N})$. Indeed, in that case, we would have
$\int_{x}^{x_{\infty}}\mathbf{D}(z,\mathbf{q})\partial_{z}\mathbf{q}(t,z)\,dz$ $=\int_{x}^{x_{\infty}}\partial_{\mathbf{q}}\mathbf{h}(z,\mathbf{q})\partial_{z}\mathbf{q}(t,z)\,dz=\int_{x}^{x_{\infty}}\frac{d}{dz}\mathbf{h}(z,\mathbf{q})-\partial_{z}\mathbf{h}(z,\mathbf{q})\,dz$ (31a)
$=\mathbf{h}(x_{\infty},\overline{\mathbf{q}})-\mathbf{h}(x,{\mathbf{q}(t,x)})-\int_{x}^{x_{\infty}}\partial_{z}\mathbf{h}(z,\mathbf{q}(t,z))\,dz.$ (31b)
Nevertheless, the integrability conditions to guarantee the existence of such a function $\mathbf{h}$ do not hold for a general system and choice of $d_{i}$’s [32]. Thus, for the application studied in Section 5 we work with a discretization of the term (30b), described in Appendix A.2.
4 The Euler equations
As an application for the techniques presented in Sections 2 and 3, we consider the one-dimensional formulation of the Euler equations for an inviscid compressible gas
$\partial_{t}\rho+\partial_{\xi}\left(\rho u\right)$ $=0,$ (32a)
$\partial_{t}\left(\rho u\right)+\partial_{\xi}\left(\rho u^{2}+p\right)$ $=0,$ (32b)
$\partial_{t}\bar{E}+\partial_{\xi}\left[(\bar{E}+p)u\right]$ $=0.$ (32c)
Here $\rho$ denotes the mass density, $u$ the flow velocity, $p$ the pressure, and $\rho u$ the momentum. The energy density given by $\bar{E}=\frac{1}{2}\rho u^{2}+\rho e$, where $\rho u^{2}$ is the kinetic energy and $e$ is the specific internal energy. Equations (32a), (32b), and (32c) describe the conservation of mass, momentum, and energy respectively, and can be derived from the Navier-Stokes equations or foundational principles in continuum mechanics if viscosity and heat flux are neglected [24].
We deal with an ideal gas with constant specific heats, hence we can define the constant heat capacity ratio $\gamma=c_{p}/c_{V}$, and our equation of state becomes [41]
$e=\frac{p}{(\gamma-1)\rho}.$ (33)
Throughout this work, we assume the density $\rho$ to be strictly positive for all times. Therefore vacuum states, which require special analytical and numerical treatment [38, Sec. 4.6], are not considered.
It is often convenient to write Equations (32) in Lagrangian form. Given a real function $\tilde{\xi}(\cdot)$, we will keep track of a set of conserved quantities at each particle defining the trajectory map [35]
[…]
From our non-vacuum state assumption $\rho$ is strictly positive, hence $x(\cdot\,,t)$ is bijective and we can use the correspondence $f(x(\xi,t),t)=f(\xi,t)$. After defining the Lagrangian specific volume
$V(x(\xi,t),t)=\frac{1}{\rho(\xi,t)},$ (35)
and the material time derivative operator (with a slight abuse of notation)
$\partial_{t}:=\partial_{t}+u\,\partial_{\xi},$ (36)
the equations (32) become
$\partial_{t}V-\partial_{x}u=0,$ (37a)
$\partial_{t}u+\partial_{x}p=0,$ (37b)
$\partial_{t}E+\partial_{x}(up)=0,$ (37c)
where $E(x,t)=\bar{E}/\rho$ [37]. Furthermore, using the equation of state (33), we have the closure relation
$p=(\gamma-1)\left(E-\frac{1}{2}u^{2}\right)/V.$ (38)
The hyperbolic system (37) can be more compactly written as
$\partial_{t}\mathbf{q}+\partial_{x}\mathbf{f}(\mathbf{q})={\bf 0},$ (39)
where $\mathbf{q}=\left(q_{1},q_{2},q_{3}\right)^{T}=\left(V,u,E\right)^{T}$ and $\mathbf{f}(\mathbf{q})=\left(q_{1},p(\mathbf{q}),q_{2}p(\mathbf{q})\right)^{T}=\left(-u,p,up\right)^{T}$ denote the vector of conserved quantities and the nonlinear flux function respectively. For smooth solutions, System (39) can be written in quasilinear form
$\partial_{t}\mathbf{q}+\mathbf{f}^{\prime}(\mathbf{q})\partial_{x}\mathbf{q}={\bf 0},$ (40)
where the Jacobian matrix
${\bf f}^{\prime}({\bf q})=\left(\begin{array}[]{ccc}0&-1&0\\ -\frac{p}{V}&\frac{{u}(1-\gamma)}{{V}}&\frac{-1+\gamma}{{V}}\\ -\frac{{p}{u}}{{V}}&p+\frac{{u}^{2}(1-\gamma)}{{V}}&\frac{{u}(-1+\gamma)}{{V}}\end{array}\right),$ (41)
has symmetric eigenvalues $\lambda_{1}(\mathbf{q})=-\sqrt{\gamma p/V}$, $\lambda_{2}=0$, $\lambda_{3}(\mathbf{q})=\sqrt{\gamma p/V}$, with corresponding eigenvectors $\mathbf{r}_{1,3}(\mathbf{q})=\left(-\frac{\sqrt{V}}{\sqrt{p}},\mp\sqrt{\gamma},\sqrt{pV}\mp u\sqrt{\gamma}\right)^{T},\,\mathbf{r}_{2}(\mathbf{q})=\left(\gamma-1,0,p\right)^{T}.$
It can be shown that the characteristic fields corresponding to the eigenvalues $\lambda_{1}$ and $\lambda_{3}$ are genuinely nonlinear while the one corresponding to $\lambda_{2}$ is linearly degenerate. Furthermore, from our non-vacuum state assumption, it follows that $p>0$ at all times, thus the system is strictly hyperbolic.
We also consider a linearization of System (39) around the state $\overline{\mathbf{q}}=(1,0,\gamma/(\gamma-1))$
$\partial_{t}\mathbf{q}+\mathbf{A}\partial_{x}\mathbf{q}={\bf 0},$ (42)
where the matrix
$\mathbf{A}=\begin{pmatrix}0&-1&0\\ -\gamma&0&\gamma-1\\ 0&\gamma&0\end{pmatrix},$ (43)
has eigenvalues $\lambda_{1}=-\gamma$, $\lambda_{2}=0$, $\lambda_{3}=\gamma$, and eigenvectors $\mathbf{r}_{1}=\left(-\frac{1}{\gamma},-1,1\right)^{T},$ $\mathbf{r}_{2}=\left(-\frac{1-\gamma}{\gamma},0,1\right)^{T},$ $\mathbf{r}_{3}=\left(-\frac{1}{\gamma},1,1\right)^{T}.$
4.1 The piston problem
Let us formulate the piston problem of gas dynamics [15], which will be the main application for the numerical treatment of boundary conditions studied in this work. The equations (32) are solved in the domain $\Omega_{t}=[\tilde{\xi}(t),+\infty[$ with a constant steady state $\overline{\mathbf{q}}=(1,0,\gamma/(\gamma-1))$ as initial condition. The left physical boundary $\tilde{\xi}(t)$ corresponds to the position of an oscillatory plate given by
[…]
where $M$ is a constant that will determine the amplitude of the oscillations. Since the particle velocity of the polytropic gas must coincide with the velocity of the plate at $\tilde{\xi}(t)$, we have the inflow boundary condition
$u(\tilde{\xi}(t),t)={\tilde{\xi}}^{\prime}(t)=-M\sin(t).$ (45)
Moreover, using the Lagrangian coordinates (34), we can solve the system of equations (39) over the time-independent domain $\Omega=[0,+\infty[$ and circumvent the difficulties arising from a moving boundary.
In the recent work [35, Ch. 3], Phan Thi My studies the treatment of outflow boundary conditions for this problem. There, the author replaces the Euler equations by a Burgers-type equation in some additional ghost cells. The simple wave behavior of the solution in this region greatly diminishes the unphysical reflections with just a few ghost cells. While using a fixed small number of ghost cells is desirable from an efficiency standpoint, it does not allow for better absorption by using a larger sponge layer.
5 Numerical results
We analyze the performance of the ABCs described in the previous sections for the nonlinear equations (39) and their linearized version (42). The setup described at the beginning of Section 2 is considered, where the inflow boundary condition $\mathbf{q}_{BC}$ is given by the piston’s motion.
5.1 Discretization details and setup
For the numerical results presented in this work, we use a high-order wave propagation formulation of Godunov-type methods, available in the Clawpack software package [29] and accessible through the Pyclaw framework [21]. For a detailed description of these methods, the reader is referred to the works [22, 20].
Denoting by $Q_{i}$ the vector of conserved quantities $\mathbf{q}$ averaged over the volume $[x_{i-1/2},x_{i+1/2}]$, our semidiscretization of Systems (39) and (42) reads
$\partial_{t}Q_{i}=-\frac{1}{\Delta x}\left(\mathcal{A}^{+}\Delta q_{i-\frac{1}{2}}+\mathcal{A}^{-}\Delta q_{i+\frac{1}{2}}+\mathcal{A}\Delta q_{i}\right).$ (46)
The fluctuations $\mathcal{A}^{\pm}\Delta q_{i-\frac{1}{2}}$ represent the net contribution of the left and right-going waves emanating from the solution of a Riemann problem at the interface $x_{i-1/2}$. The initial data for each Riemann problem is given by a piecewise polynomial reconstruction of the solution at both sides of the interface. To avoid spurious oscillations around strong shocks without using an extremely refined grid, a piecewise linear reconstruction with the TVD minmod limiter is used. An advantage of the semi-discretization (46) is that it can be applied to systems in nonconservative form, like (3), in a natural way [2, 22]. Since the semidiscretization and numerical inflow boundary condition (see Equations (48) and (49) below) formally attain second order of accuracy, we use Heun’s two-stage second-order accurate Runge-Kutta scheme to integrate System (46) in time.
To compute an approximate Riemann solution at an interface with left and right constant initial data $q_{L},\,q_{R}$, we use an HLL Riemann solver [11]. The approximate speeds are defined to bound the wave speeds from the exact Riemann solution from below and above,
$s^{1}=\min_{p}\left(\min\left(\lambda_{p}^{L},\lambda^{R}_{p}\right)\right)=\min\left(\lambda_{1}^{L},\lambda_{1}^{R}\right),\quad s^{2}=\max_{p}\left(\max\left(\lambda_{p}^{L},\lambda^{R}_{p}\right)\right)=\max\left(\lambda_{3}^{L},\lambda^{R}_{3}\right).$ (47)
Since the eigenvalues of our Jacobian matrix are symmetric, we have $s^{1}=-s^{2}$. Therefore, in this case, the HLL solver coincides with the Local Lax-Friedrichs (or Rusanov) solver.
The computation of fluctuations and the reconstruction process require the use of ghost cells at the left and right boundaries of the domain, denoted by $\mathcal{C}_{-1}$ and $\mathcal{C}_{N_{x}}$ respectively. We assume the piston (of infinitesimal width) to be located at the interface between the cells $\mathcal{C}_{-1}$ and $\mathcal{C}_{0}$ ($x=0$). Using the flow conditions at the piston, formulated in [7], and their discretization described in [35], we get the second-order accurate approximations
$u^{[n]}_{-1}=-2M\sin(t^{[n]})-u_{0}^{[n]},\quad p^{[n]}_{-1}=p_{0}^{[n]}-M\cos(t^{[n]})\Delta x,$ (48a)
$V^{[n]}_{-1}=V^{[n]}_{0}\frac{\gamma+\delta_{p}}{\gamma-\delta_{p}},\quad\delta_{p}=\frac{p^{[n]}_{0}-p^{[n]}_{-1}}{p^{[n]}_{0}+p^{[n]}_{-1}},\quad E_{-1}^{[n]}=\frac{1}{\gamma-1}p_{-1}^{[n]}V_{-1}^{[n]}+\frac{1}{2}(u_{-1}^{[n]})^{2}.$ (48b)
Following an analogous derivation for the linearized equations (42), we obtain
$u^{[n]}_{-1}=-2M\sin(t^{[n]})-u^{[n]}_{0},\quad p^{[n]}_{-1}=p^{[n]}_{0}-M\cos(t^{[n]})\Delta x,$ (49a)
$V^{[n]}_{-1}=\frac{1}{\gamma}(p^{[n]}_{0}-p^{[n]}_{-1}+\gamma V^{[n]}_{0}),\quad E^{[n]}_{-1}=\frac{1}{\gamma-1}(p^{[n]}_{-1}+\gamma V^{[n]}_{-1}),$ (49b)
where we have the closure relation $p=\gamma V+(\gamma-1)E$ instead of (38). At the right boundary we set $Q_{N_{x}}=\overline{\mathbf{q}}$, where $\overline{\mathbf{q}}$ is the far-field state $\overline{\mathbf{q}}=(1,0,\gamma/(\gamma-1))$ for the nonlinear equations (39), and $\overline{\mathbf{q}}=\bf 0$ for the linear equations (42). Here, $\mathcal{C}_{N_{x}}$ is the ghost cell centered at $x_{\infty}$. For both systems, the initial condition will be their corresponding far-field state $\overline{\mathbf{q}}$.
For all the numerical tests below, $x_{s}=20\pi$, and $x_{\infty}$ takes different values to evaluate the dependence of the ABCs on the sponge layer’s length $\omega=|x_{\infty}-x_{s}|$. Since the piston’s frequency is $(2\pi)^{-1}$, as pointed out in Equation (44), the characteristic wavelength of the solution will be $L=2\pi$. Different spatial refinement levels are considered, such that $N$ cells are contained in a wavelength $L$, for $N\in\{10,50,250\}$. A Courant number $C=0.8$ is used to compute $\Delta t$ at each time step under the standard CFL condition $\Delta t={C\Delta x}/{\max_{i}\max_{p}{|\lambda_{p}(Q_{i})|}}.$ The amplitude of the piston’s oscillations is set to $M=0.4$, which in the nonlinear case will lead to a solution dominated by strong shocks, as studied by Inoue and Yano [15].
Given a refinement level $N$ and a sponge layer length $\omega$, we define the error associated with an ABC as
$E_{ABC}(N,\omega)=\max_{t\in[0,40\pi]}\frac{{\|u^{[N]}_{\text{ref}}(t)-u^{[N,\omega]}_{\text{ABC}}(t)\|}_{L^{1}(0,x_{s})}}{\|u^{[N]}_{\text{ref}}(t)\|_{L^{1}(0,x_{s})}},$ (50)
where we use a grid function norm over the cell-averaged velocity $u$, and the reference solution is computed on a larger computational domain such that no wave interaction takes place at the boundary. Similarly, the error corresponding to the numerical discretization is approximated by
$E_{num}(N)=\max_{t\in[0,40\pi]}\frac{{\|u^{[N]}_{\text{ref}}(t)-u^{[3000]}_{\text{ref}}(t)\|}_{L^{1}(0,x_{s})}}{\|u^{[N]}_{\text{ref}}(t)\|_{L^{1}(0,x_{s})}},$ (51)
where $u^{[3000]}_{\text{ref}}(\cdot)$ is appropriately undersampled.
5.2 Parameter configuration
Before proceeding with a performance comparison between the different ABCs, we tune the following parameters:
  • a slowing-down ($s$) and a damping ($d$) function for the far-field operators-based techniques (SDO, S-SDO, NDO),
  • a scalar weight function ($\Gamma$) for the relaxation methods (RM, RM-M).
These functions are constructed using $\Gamma_{A}$ (9) and $\Gamma_{B}$ (with $b=1/2$) (11), by taking
$\Gamma(x)=\Gamma_{(\cdot)}(x),\quad s_{(\cdot)}(x)=\Gamma_{(\cdot)},\quad d_{(\cdot)}(x)=\sigma\left(1-\Gamma_{(\cdot)}(x)\right),$ (52)
where $\sigma$ is the maximum damping rate attained in the sponge layer. The shapes of these functions are illustrated in Figure 3.
Figure 2: Weight functions $\Gamma_{(\cdot)}$ and damping functions $d_{(\cdot)}$, as defined in Equations (9), (11), and (52) respectively
Figure 3: Relative error (50) for the nonlinear equations (39) with SDO ABC and different maximum damping rates $\sigma$ $(N=250)$
First, we evaluate different combinations of $s(\cdot)$ and $d(\cdot)$ for SDO with $\sigma=30$, $N=250$, and considering different sponge layer lengths $\omega$. The relative errors obtained with each combination are displayed in Table 1. For the linear equations (42), while the choice of $s$ and $d$ does not seem to be critical, the best performance is attained with $s_{A}$ and $d_{B}$. With this configuration, the slowing-down effect is greater than the damping at the early stage of the absorbing process. Thus, the length of the outgoing waves is reduced before they are damped and, consequently, the drawbacks associated with large wavelengths, addressed in Section 2, are alleviated.
{tblr}
cells = c, column3 = Mercury, column8 = Mercury, cell11 = r=2, cell12 = c=4, cell16 = c=4, vlines, hline1,3,7 = -, hline2 = 2-9, rowsep = 1pt, $\frac{\omega}{L}$ Linear equations (42) Nonlinear equations (39)
$s_{A}$, $d_{A}$ $s_{A}$, $d_{B}$ $s_{B}$, $d_{A}$ $s_{B}$, $d_{B}$ $s_{A}$, $d_{A}$ $s_{A}$, $d_{B}$ $s_{B}$, $d_{A}$ $s_{B}$, $d_{B}$
$1/8$ 4.53e-04 4.46e-03 1.08e-03 2.74e-03 6.54e-04 5.63e-02 2.50e-04 2.66e-03
$1/4$ 4.46e-04 5.29e-05 6.92e-04 3.62e-04 2.52e-01 6.87e-01 3.62e-04 4.53e-04
$1/2$ 3.33e-04 5.11e-05 4.79e-04 1.66e-04 3.73e-01 3.86e-01 4.05e-04 3.96e-04
$1$ 2.03e-04 5.69e-05 2.44e-04 6.06e-05 4.10e-04 4.82e-01 4.13e-04 3.61e-04
Table 1: Error (50) obtained with SDO and different slowing-down and damping functions $(\sigma=30,\,N=250)$
In contrast with the linear scenario, the choice of slowing-down and damping functions drastically affects the behavior of the SDO ABC for the nonlinear equations (39) (see Table 1). For this system, $s_{A}$ and $d_{B}$ yield the least error overall. This can be attributed to the slowing-down operator’s tendency to generate entropy inside the sponge layer, illustrated in Appendix A.3, which needs to be dissipated by a rapidly increasing damping rate. The effect of using different maximum damping rates $\sigma$ with SDO is depicted in Figure 3. Since no significant improvement is obtained for larger values, $\sigma=30$ is used for the SDO ABC in the following. Proceeding analogously, we select $d_{B}$ and $\sigma=20$ for the NDO ABC, as suggested by the results displayed in Table 5 and Figure 5.
Finally, the scalar weight function $\Gamma$ for the RM and RM-M techniques is chosen similarly. As can be seen in Table 2, the best performance is obtained with $\Gamma_{B}$. Nevertheless, these techniques do not seem to be critically dependent on the choice of $\Gamma$, even in the nonlinear setting.
{tblr}
cells = c, column3 = Mercury, cell11 = r=2, cell12 = c=2, vlines, hline1,3,7 = -, hline2 = 2-3, rowsep = 1pt, $\frac{\omega}{L}$ Nonlinear equations (39)
$d_{A}$ $d_{B}$
$1/8$ 8.93e-03 5.74e-03
$1/4$ 5.10e-03 3.48e-03
$1/2$ 3.74e-03 2.80e-03
$1$ 3.21e-03 2.03e-03
Figure 4: Relative error (50) for the nonlinear equations (39) with NDO ABC 3.4 and different damping functions $(\sigma=20,\,N=250)$
Figure 5: Relative error (50) for the nonlinear equations (39) with NDO ABC and different maximum damping rates $\sigma$ $(N=250)$
{tblr}
cells = c, column3 = Mercury, column5 = Mercury, column7 = Mercury, column9 = Mercury, cell11 = r=3, cell12 = c=4, cell16 = c=4, cell22 = c=2, cell24 = c=2, cell26 = c=2, cell28 = c=2, vlines, hline1,4,8 = -, hline2-3 = 2-9, rowsep = 1pt, $\frac{\omega}{L}$ Linear equations (42) Nonlinear equations (39)
RM RM-M RM RM-M
$\Gamma_{A}$ $\Gamma_{B}$ $\Gamma_{A}$ $\Gamma_{B}$ $\Gamma_{A}$ $\Gamma_{B}$ $\Gamma_{A}$ $\Gamma_{B}$
$1/8$ 3.30e-03 2.16e-03 8.64e-04 1.36e-03 8.06e-04 9.69e-04 1.24e-03 5.18e-04
$1/4$ 1.94e-03 7.75e-04 6.54e-04 5.04e-04 1.06e-03 1.09e-03 3.78e-04 2.42e-04
$1/2$ 8.78e-04 2.31e-04 4.40e-04 2.23e-04 1.10e-03 2.13e-03 2.02e-03 4.61e-04
$1$ 3.05e-04 6.76e-05 3.05e-04 8.71e-05 1.09e-03 9.81e-04 4.37e-04 4.13e-04
Table 2: Error (50) obtained with the scalar 2.2 and matrix-valued 3.2 RMs and different weight functions $(N=250)$
5.3 Performance comparison
Let us compare the behavior of the different relaxation methods discussed in Sections 2.2, 3.2, and 3.3. We refer to Algorithms 1 and 2 as RM and RM-M respectively, and to their use described in Algorithm 3 as RM 2 and RM-M 2. For the latter, the weight function $\Gamma_{B}^{2}$ is used, according to the observations from Section 3.3. Moreover, we consider the application of these techniques after each stage of the Runge-Kutta time integration, as proposed by Wasistho et al. [39], and refer to them as RM RK and RM-M RK. The results are displayed in Figures 7 and 7 for the linear (42) and nonlinear (39) equations respectively, along with the errors corresponding to the numerical discretization (51) and the use of constant extrapolation at the right boundary.
For the linear equations, there is no perceivable difference between RM and RM-M, and their variants. This can be attributed to the fact that due to the piston motion, all the waves that enter the sponge layer are right-going, and no interaction between them takes place in the linear setting. Therefore, there is no difference betwen absorbing waves selectively or not. Conversely, for the nonlinear equations, there is a perceivable improvement with the RM-M with respect to the RM when larger portions of a characteristic wavelength $L$ are contained in the sponge layer. The techniques RM and RM-M attain exactly the same error as RM 2 and RM-M 2 respectively for both systems. This validates the equivalence between Algorithms 2 and 3, established in Section 3.3 for the linear equations, and suggests that it extends to the nonlinear case. Additionally, as the grid is refined, we see that the error obtained with RM-M-RK decays in a smoother than the other relaxation methods as the sponge layer length $\omega$ increases.
Figure 6: Comparison of relaxation methods for the linear equations (42)
Figure 7: Comparison of relaxation methods for the nonlinear equations (39)
Figure 8: Comparison of different ABCs for the linear equations (42)
Figure 9: Comparison of different ABCs for the nonlinear equations (39)
Now we assess the perfomance of all the ABCs addressed in Sections 2 and 3; the results are displayed in Figures 9 and 9. For the linear equations, the behavior of the different ABCs is similar for all the grid sizes, with the exception of S-SDO, which attains a larger error that diminishes as the grid is refined.
For the nonlinear equations, the methods based on far-field operators (SDO, S-SDO, NDO) fail to provide a satisfactory error and the numerical solution diverges for the coarsest grids. This suggests that the discretization is not stable with the time step size prescribed by the CFL condition with the chosen Courant number. Indeed, we can see that the $L^{2}$ norm of $\mathbf{D}(x)$ in Equation (3) is $\max_{i}|d_{i}s_{i}\lambda_{i}|$, which is larger than time step size restriction for a coarse grid when the maximum damping parameter $\sigma$ is large. As seen in Figure 3, a large maximum damping rate might be necessary to attain a prescribed error for a fixed sponge layer length. Since it is convenient to use an explicit scheme for the hyperbolic part of the problem, a way to deal with this artificial stiffness, not explored in this work, is to use an IMEX scheme [33] to evolve the semidiscretization. On the other hand, the RM and its extension RM-M, introduced herein, show to be the most robust ABCs for both the linear and nonlinear equations.
Once the grid (and thus the time step size) is refined, and the stability issues are overcome, we can see that the SDO ABC attains its minimum error for a sponge layer length of around $0.08L$. It is followed by RM-M and NDO which attain their minimum error for sponge layer lengths of $0.6L$ and $5L$ respectively. While the RM attains a low error for a sponge layer length of $0.04L$, it still does not reach the absborption of the other methods, even for a sponge layer length of $10L$. In our particular setting, we know that the first two shocks of the wave train are periodically merged in the far-field [15], and such interactions might cause information to propagate back into the computational domain [12]. Consequently, as the sponge layer size increases, methods that selectively damp a characteristic field, like SDO, NDO, and RM-M, allow this physical information to recirculate, while scalar techniques, like RM and S-SDO, fail to capture this behavior (see Remark 1 below).
Finally an execution time comparison between the different ABCs for the nonlinear equations (39) is displayed in Table 3. We consider the best wall clock time obtained for each method after 7 runs on a MacBook Pro 2019 with a 2.3 GHz 8-Core Intel Core i9 processor. As expected, the scalar methods, RM and S-SDO, are the fastest ones, followed by the directional methods RM-M SDO. The NDO technique is, by a large margin, the least efficient due to the computationally expensive treatment of its nonlocal integral term (see Appendix A.2).
Remark 1.
To further illustrate the effect of damping waves corresponding to a particular characteristic field, we consider an artificial setup where an incoming pulse is created inside the sponge layer as shown in Figure 10. The initial pulse is immediately damped by the scalar methods (RM and S-SDO), while just the right-going components of it are absorbed by the directional methods (RM-M and SDO).
Figure 10: Above: Effect of directional (SDO) and scalar (SSDO) damping operators applied to an incoming physical wave. The sponge layer is colored light grey. Below: Effect of directional relaxation method with matrix-valued weight function (RM-M) and scalar relaxation method (RM) applied to an incoming physical wave.
N RM RM-M S-SDO SDO NDO
10 0.356278 0.413172 0.311402 0.338358 0.487306
50 1.37656 1.89572 1.49388 1.32294 4.84846
250 14.7071 31.6987 10.8559 32.3652 107.529
Table 3: Execution time comparison of ABCs with different refinement levels $N$ for the nonlinear equations with a sponge layer of length $\omega=10L$
6 Conclusion
Methods from two families of absorbing boundary conditions were studied, and their performance was tested for the piston problem of gas dynamics. The relaxation method with a scalar-valued weight function is the least expensive and simplest to implement. Karni’s slowing-down and damping operators are able to selectively absorb waves from a specific characteristic field, allowing physical waves to circulate back into the computational domain. However, this technique requires the addition of an artificial source term that might become stiff depending on the prescribed damping rate, leading to additional time step size restrictions to maintain stability with an explicit time integrator. Among the extensions for both techniques presented in this work, the relaxation method with a matrix-valued weight function is an attractive alternative, as it preserves the remarkable robustness of the relaxation method (even in the nonlinear setting), while selectively damping outgoing waves. We expect this property to be more advantageous in a case where more interactions take place inside the sponge layer, for instance, when the sponge layer is used as a wavemaker (i.e. nonconstant far-field state), which is a common application for the relaxation method. Furthermore, sophisticated scalar weight functions (e.g. that depend on the conserved quantities), that have already been developed for the relaxation method, can be used with this extension in a straightforward manner. Despite being the least efficient technique considered in this work, we expect the nonlinear damping operator to be more effective for high-order discretizations, where a high-order approximation of the solution is available inside each cell. Thus, more information can be used for the computation of the integral damping term.
Acknowledgements
I would like to express my deepest gratitude to Prof. David Ketcheson and Prof. Giovanni Russo for their invaluable advice and support throughout the development of this work.
Appendix A Appendix
A.1 Construction of damping operators
The construction of the damping operators and the matrix-valued weight function of Sections 2.1, 3.4, and 3.2, relies on the computation of a matrix $\mathbf{M}=\mathbf{R}\mathbf{\Sigma}\mathbf{R}^{-1}$, where $\mathbf{\Sigma}=\text{diag}(\sigma_{i})$, and $\mathbf{R}$ contains the right eigenvectors of the flux Jacobian matrix as columns. For the linear equations (42) we have
$\mathbf{M}=\left(\begin{array}[]{ccc}\frac{1}{\gamma}\left(\frac{p_{1}+\sigma_{3}}{2}+(\gamma-1)\sigma_{2}\right)&\frac{1}{2\gamma}(\sigma_{1}-\sigma_{3})&\frac{\gamma-1}{\gamma^{2}}\left(\frac{-\sigma_{1}-\sigma_{3}}{2}+\sigma_{2}\right)\\ \frac{1}{2}(\sigma_{1}-\sigma_{3})&\frac{1}{2}(\sigma_{1}+\sigma_{3})&\frac{\gamma-1}{2\gamma}(\sigma_{3}-\sigma_{1})\\ -\frac{1}{2}(\sigma_{1}+\sigma_{3})+\sigma_{2}&\frac{1}{2}(\sigma_{3}-\sigma_{1})&\frac{1}{\gamma}\left(\frac{\gamma-1}{2}(\sigma_{1}+\sigma_{3})+\sigma_{2}\right)\end{array}\right),$ (53)
while for the nonlinear equations (39) we get
$\mathbf{M}=\begin{pmatrix}\frac{1}{2\gamma}(2\gamma\sigma_{2}+C_{1})&\frac{1}{2p\gamma}(\sqrt{\gamma pV}C_{2}+uC_{3}C_{1})&\frac{1}{2p\gamma}(-C_{1}C_{3})\\ \frac{C_{2}p}{C_{4}}&\frac{\sqrt{\gamma}\sqrt{pV}(\sigma_{1}+\sigma_{3})+uC_{3}C_{2}}{C_{4}}&\frac{-C_{2}C_{3}}{C_{4}}\\ \frac{-pVC_{1}+C_{2}\sqrt{pV}\sqrt{\gamma}u}{2V\gamma}&\frac{\sqrt{pV}uC_{1}+\sqrt{\gamma}(C_{3}u^{2}-pV)C_{2}}{C_{4}\sqrt{\gamma}}&\frac{-\sqrt{pV}C_{1}-\sqrt{\gamma}C_{3}C_{2}u+\gamma\sqrt{pV}(\sigma_{1}+\sigma_{3})}{C_{4}\sqrt{\gamma}}\end{pmatrix},$ (54)
where $C_{1}=\sigma_{1}-2\sigma_{2}+\sigma_{3}$, $C_{2}=\sigma_{1}-\sigma_{3}$, $C_{3}=\gamma-1$, and $C_{4}=\sqrt{\gamma pV}$.
A.2 Discretization of integral source term with nonconservative product
Let us consider the PDE
$\partial_{t}\mathbf{q}(t,x)+\partial_{x}\mathbf{f}(\mathbf{q}(t,x))=\int_{x}^{x_{\infty}}\mathbf{M}(z,\mathbf{q})\partial_{z}\mathbf{q}(t,z)\,dz,$ (55)
under the Ansatz that the solution is piecewise affine inside the cells $\mathcal{C}_{i}$ and not necessarily continuous at the interfaces, e.g.,
$\mathbf{q}(x,t)=\sum_{i}^{m}\chi_{\mathcal{C}_{i}}(x)\left[Q_{i}(t)+(x-x_{i})\sigma_{i}(t)\right],$ (56)
where $x_{i}$ is the center of the cell $\mathcal{C}_{i}$, $x_{\infty}$ the right interface of $\mathcal{C}_{m}$, and $\sigma_{i}(t)$ is a vector of slopes.
We can write (55) at a point $x\in\mathcal{C}_{j}$ ($x<x_{j}$ for simplicity) as
$\partial_{t}\mathbf{q}+\partial_{x}\mathbf{f}(\mathbf{q})$ $=\int_{x}^{x_{j}}\mathbf{M}\partial_{z}\mathbf{q}\,dz+\sum_{i=j}^{m-1}\int_{x_{i}}^{x_{i+1/2}^{-}}\mathbf{M}\partial_{z}\mathbf{q}\,dz+\int_{x_{i+1/2}^{+}}^{x_{i+1}}\mathbf{M}\partial_{z}\mathbf{q}\,dz+\int_{\{x_{i+1/2}\}}\mathbf{M}\partial_{z}\mathbf{q}\,dz,$ (57)
[…]
where ${x_{i+1/2}^{-}}$ $x_{i+1/2}^{+}$ represent the left and right limits of the interface point $x_{i+1/2}$. Simpson’s rule is used to approximate the interior integrals in Equation (58). Since $\mathbf{q}$ is discontinuous at $x_{i+1/2}$, we interpret the nonconservative product as a Borel measure $\mu$ such that
[…]
where $\phi$ is a Lipschitz continuous function connecting $\mathbf{q}^{-}(t)=\lim_{x\to x_{i+1/2}^{-}}\mathbf{q}(x,t)$ with $\mathbf{q}^{+}(t)=\lim_{x\to x_{i+1/2}^{+}}\mathbf{q}(x,t)$ [26]. Here we choose $\phi$ to be simply a straight line $\phi(s)=s\mathbf{q}^{+}+(1-s)\mathbf{q}^{-}$, and approximate the integral in (59) with the midpoint rule.
A.3 Stability of the slowing-down operator for nonlinear equations
One can show that the slowing-down operator does not destroy the $L^{2}$ stability of a linear system. Indeed, we can consider a general linear system of $N$ hyperbolic PDEs
$\partial_{t}\mathbf{q}+\mathbf{A}\partial_{x}\mathbf{q}=\bf 0,$ (60)
defined over the real line such that $\lim_{x\to\pm\infty}\mathbf{q}(t,x)=\bf 0$. It is then clear that $\partial_{t}||\mathbf{q}(t,\cdot)||_{L^{2}()}=0$. Let us analyze the effect of the slowing-down operator applied over the sponge layer $[x_{s},x_{\infty}]$. For instance, we absorb the right-going waves, i.e., we solve
$\partial_{t}\mathbf{q}+\mathbf{S}(x)\partial_{x}\mathbf{q}=\bf 0,$ (61)
where $\mathbf{S}(x)=\mathbf{R}\mathbf{\Lambda}_{\mathbf{S}}(x)\mathbf{R}^{-1}$ such that $\mathbf{A}=\mathbf{R}\mathbf{\Lambda}\mathbf{R}^{-1}$, and the main entries of the diagonal matrix $\mathbf{\Lambda_{S}}(x)$ are given by
$\tilde{\lambda}_{i}(x)=\begin{cases}\lambda_{i},\quad&\text{if }\lambda_{i}\leq 0,\\ s(x)\lambda_{i},\quad&\text{if }\lambda_{i}>0.\end{cases}$ (62)
The scalar smooth slowing-down function $s(x)$ strictly decreases from $1$ to 00 inside the sponge layer $[x_{s},x_{\infty}]$ and is constant everywhere else. Then, multiplying (62) by $\mathbf{q}^{T}$ and integrating in space, we get
[…]
The inequality is a consequence of $\mathbf{S}^{\prime}(x)$ being semi-negative definite, since ${\tilde{\lambda}_{i}}^{\prime}(x)\leq 0$.
The same cannot be said in the nonlinear setting, even for a scalar PDE. For instance, we can apply the slowing-down operator, described above, to Burgers’ equation
$\partial_{t}u+\frac{1}{2}\partial_{x}u^{2}=0,$ (64)
for which we know that $\partial_{t}||u||_{L^{2}()}\leq 0$ holds. The modified equation reads
$\partial_{t}u+s(x)\frac{1}{2}\partial_{x}u^{2}=0,$ (65)
and after some manipulations, it yields
$\frac{1}{2}\partial_{t}\int_{u}^{2}\,dx=-\frac{1}{3}\int_{s}(x)\partial_{x}u^{3}\,dx=-\frac{1}{3}\int_{\partial}{}_{x}[s(x)u^{3}]-s^{\prime}(x)u^{3}\,dx=\frac{1}{3}\int_{x_{s}}^{x_{\infty}}s^{\prime}(x)u^{3}\,dx.$ (66)
If $u<0$ inside the sponge layer, the last term (thus $\partial_{t}||u||_{L^{2}()}$) becomes strictly greater than 00 and the entropy-dissipative nature of the original system is lost.